OpenStudy (anonymous):

sin theta= 3/5 cos beta = 2sqrt6/5 Find cos(theta+beta). Is the answer 8sqrt6-3/25

OpenStudy (anonymous):

\[\cos \left( \theta+\beta \right)=\cos \theta \cos \beta -\sin \theta \sin \beta \] find \[\cos \theta ~and~ \sin \beta \] then solve

OpenStudy (anonymous):

Yes I need someone to check my answer

OpenStudy (anonymous):

show me I will check